# QPET: A Versatile and Portable Quantity-of-Interest-Preservation Framework for Error-Bounded Lossy Compression

### v3

@@ 6.2.1. Outperforming speeds
First, we measure and profile the execution speeds of baselines and QPET-integrated compressors to examine how QPET has accelerated the QoI-preserving error-bounded lossy compression. From the compression and decompression speeds on the Anvil cluster with the QoI functions $Q(x)=x^{2}$ and $Q(x)=x^{3}$ and the relative QoI error threshold $\tau=1\mathrm{e}{-3}$, reported in Figure 3 to 4 (no results from QoI-SZ3/HPEZ for $Q(x)=x^{3}$ because their current implementation don’t support this QoI), we can conclude as follows: compared to existing parameter-search-based baselines and QoI-preseving compressors, QPET-integrated compressors have significantly improved the compression throughputs (4x $\sim$ 6x of *-OptZ-R and 2x $\sim$ 2.5x of QoI-SZ3/HPEZ), with similar decompression throughput with the fast general-purpose compressors. Parameter search on general-purpose compressors (*-OptZ-R) need to perform repeated executions for acquiring a certain QoI error threshold, and in our evaluations, this task may cost 4 to 10 or even more compression executions and QoI validations. Moreover, QPET brings quite limited time overheads in the decompression process (because neither error-bound computation nor QoI validation is needed), and sometimes it even improves the decompression speed over baselines due to the reduced accuracy required in the compression to ensure the same QoI error threshold. Last, Compared to the existing QoI-SZ3/HPEZ, for the same use cases, QPET gains 100% $\sim$ 150% compression speed improvement and up to 400% decompression speed improvement, which is highly attributed to the efficient error-bound selection and tuning process proposed by QPET, and also the revised implementation in storing and recovering the point-wise error bounds.
@@ 6.2.2. Representative showcases
Table 5. Showcases of QoI-preserving errorlossy compression. For blocked QoIs, $n=4^{3}$ (i.e., average on 4x4x4 blocks). $\epsilon$: data error bound. $\tau$: QoI error threshold. #It: Number of iterations. CR: Compression ratio. $T_{c}$: Compression throughput (in MB/s). $T_{d}$: Decompression throughput (in MB/s). All experiments are performed on Purdue Anvil, and all error bounds are relative (absolute value/value range). N/A indicates that the baseline is not available in the case.
@@
From those showcases, key findings are: (1) Matching the discussion in Section 6.2.1, QPET-integrated compressors deliver better compression throughputs than all baselines. The QPET-integrated compressor achieves 2x to 10x compression speedups over the parameter-search-based solutions. Moreover, SZ3/HPEZ-QPET has 2x to 3x compression speeds and 3x to 6x decompression speeds over QoI-SZ3/HPEZ. (2) With better throughputs, QPET-integrated compressors bring no worse (and mostly improved) compression ratios over the baselines. On $q(x)=x^{2}$, under error bounds of $10^{-4}$, QPET gained similar compression ratios with baselines in much shorter time costs. On preserving $q(x)=x^{3}$ of RTM-3500 and Hurricane-Cloud data (they both feature small absolute values less than 0.002), QPET-integrated compressors achieved 30% - 700% compression ratio improvements over the best parameter-search solution. On NYX-Velocity_x data ranging from zero to tens of millions, general-purpose compressors failed to preserve the QoI $\tanh{x}$ well due to the required point-wise data accuracy extremely diverging among small and large values. In contrast, QPET handled this case excellently. On multivariate QoIs, QPET significantly improved the compression ratio over baselines, as benefited from Theorem 5.4. Compared with existing QoI-preserving compressors on basic QoIs, QPET-integrated compressors still achieve noticeable compression ratio gains. On preserving $q(x)=\log_{2}x$ on Scale-LetKF-T data, SPERR-QPET achieved 100% CR improvement over QoI-HPEZ under $\epsilon=\tau=10^{-4}$. On preserving $q(X)=\frac{1}{n}\sum x_{i}^{2}$ on SegSalt-Pressure-3000 data, SPERR-QPET achieved 133% CR improvement over QoI-HPEZ under $\tau=10^{-3}$. General-purpose compressors usually over-preserve data quality to preserve the QoI error threshold because they can neither apply point-wise data accuracy nor auto-correct outliers. This explains why non-QoI-integrated compressors have limited compression ratios on QoI-preserving compression tasks, even after parameter searches on optimizing the compression ratio. Meanwhile. Existing QoI-preserving compressors (like QoI-SZ3/HPEZ and MGARD-QoI) bring significant computational overheads for the QoI-preserving compression tasks. With lightweight error-bound tuning and outlier correction mechanisms, QPET successfully avoided the abovementioned limitations, proving its effectiveness and usability in preserving the QoI.
@@ 6.2.3. Improved Rate-distortions
To comprehensively understand the compression ratio improvement brought by QPET in the QoI-preserving error-bounded data compression, the compressors mentioned above are evaluated on a large variety of QoI functions and error thresholds. Then, we plot the bit rates and maximum QoI errors (relative, absolute value divided by value range), showing that QPET-integrated compressors have substantially improved the compression ratio when achieving the same maximum QoI errors as others. Due to the page limit, we will show the most representative results among the excellent outcomes we acquired.
@@
Regarding different point-wise QoI functions, Figure 5 presents the bit rate-maximum QoI error plots we acquired on 6 data domains. In this figure and all other similar ones in the paper, the results from the same base compressors are plotted in the curves of the same color, with QPET-integrated compressors solid and baselines dashed. Moreover, the results from QoI-SZ3/HPEZ appear if and only if the corresponding) QoI is supported by them. In Figure 5, QPET-integrated compressors have shown the optimized compression ratio (bit rate) in every test case and achieve substantial improvements in most. For example, on the Scale-LetKF dataset (Figure 5 (b)), SPERR-QPET has around 100%/50% compression ratio improvements over the best baseline (QoI-SZ3/HPEZ) when the error threshold for the square of data ($Q(x)=x^{2}$) is set as $1\mathrm{e}{-2}$/$1\mathrm{e}{-3}$ (relative, and all the followings are same). On other QoIs for which existing QoI-preserving data compressors fail to support, such as $x^{3}$ and $\tanh{x}$ (Figure 5 (c, d, f)), QPET makes great use of the point-wise data accuracy to preserve QoIs, significantly outperforming parameter-search-based baselines (*-OptZ-R) by up to over 1000% compression ratio improvement (on NYX dataset with $Q(x)=x^{3}$) and higher throughputs. Regarding preserving $Q(x)=\sin{10x}$, which is close to a linear function in several local data regions, point-wise data accuracy brings a limited contribution to the compression, so the compression ratios of QPET-integrated compressors are not outperforming parameter-search-based baselines. Nevertheless, QPET-integrated compressors can acquire the same compression ratios at much lower time costs, which is still a critical improvement for high-performance data processing tasks.
@@
Multivariate QoI functions contain two sub-categories: QoI on data blocks within one data field and QoI vector data among multiple fields. In Figure 6, for better comparison with baselines (particularly QoI-SZ3/HPEZ and MGARD-QoI), we perform evaluations of QPET and the baselines on the average of $x$ (data value itself), $x^{2}$ (square of the data value), and $x^{3}$ (cubic of the data value) on partitioned fixed-size data blocks. On Figure 6 (a) and (b), we can find that MGARD-QoI and QoI-SZ3 present unstable and suboptimal compression ratios, and QoI-HPEZ cannot outperform HPEZ-QPET and SPERR-QPET with an obvious gap between them. For $Q(X)=\frac{1}{n_{b}}\sum x^{2}$ and $Q(X)=\frac{1}{n_{b}}\sum x^{3}$, the 3 QPET-integrated compressors are the top-3 of all compressors, outperforming both parameter-search-based baselines and QoI-SZ3/HPEZ. On preserving the $Q(X)=\frac{1}{n_{b}}\sum x^{2}$ when compressing the RTM dataset (Figure 6 (c)), SPERR-QPET gains an aggregated compression ratio of $\approx$ 215 (bit rate of $\approx$0.15) under the QoI error tolerance of $1\mathrm{e}{-4}$, which is 108% better than the best baseline SPERR-OptZ-R (CR $\approx$ 103). On preserving the $Q(X)=\frac{1}{n_{b}}\sum x^{3}$ when compressing the Hurricane dataset (Figure 6 (e)), HPEZ-QPET gains aggregated compression ratios of $\approx$ 121/45.4 under the QoI error tolerance of $1\mathrm{e}{-3}$/$1\mathrm{e}{-4}$, which is 200%/220% of HPEZ-OptZ-R compression ratios ( $\approx$ 60.9/20.6). The improvement of compression ratios in those test cases is highly attributed to Theorem 5.4 because of the large number of variables (same as the block size 64) in the QoIs so that QPET can apply much higher point-wise QoI ($x$/$x^{2}$/$x^{3}$) error thresholds on each data point compared to the overall error threshold.
@@ 6.2.4. Ablation study
At the end of our evaluation, we commit an ablation study of QPET design components, separately validating how different strategies described in Algorithm 2,  3, and  4 contribute to the performance of QPET.
On bounding the error of $Q(x)=x^{3}$, Figure 8 (a) and (b) compare QPET to its derivation QPET-A, which uses the analytical solution of inequality $\left|(\epsilon_{i}+x_{i})^{3}-x_{i}^{3}\right|\leq\tau$ instead of Algorithm 2 to determine the point-wise error bound $\epsilon_{0}$ for $x_{0}$. On both of the datasets, QPET achieves the same compression ratio as QPET-A, meaning that Algorithm 2 performed well to replace analytic solutions. This fact indicates that the mechanism of QPET has worked well in determining the compression error bounds.
Next, we investigate the impact of Throrem 5.4 and the parameter $c$ by evaluating SPERR-QPET on the SCALE-LetKF with the QoI of the average of $x^{2}$ on 4$\times$4$\times$4 data blocks with different configurations. Under $c=0$, SPERR-QPET deactivates Theorem 5.4, and larger $c$ brings large error bound estimations from Throrem 5.4. Figure 8 (c) illustrates the corresponding results. Higher $c$ does improve compression by proposing higher and still trustworthy estimations of point-wise error bounds. As the improvement of compression ratios becomes minor when $c$ arrives at 3.0, we selected this value as the default in SPERR-QPET.
@@ 7. Conclusion
In this work, we propose QPET, a versatile and portable framework for error-bounded lossy compression that enables the preservation of a broad range of Quantity-of-Interest (QoI) while significantly improving the performance and quality of existing QoI-preserving error-bounded lossy compression. We further integrate QPET into 3 state-of-the-art error-bounded lossy compressors of different archetypes. Experimental results demonstrate that QPET is faster than all existing QoI-preserving error-bounded lossy compression solutions, leading 2x to 10x compression speedups over the baselines of both existing parameter-search methods and QoI-preserving compressors. QPET also delivers up to 1000% compression ratio improvements over existing QoI-preserving error-bounded lossy compression solutions. In future research, we will optimize QPET in the following three aspects: (1) Improve the inclusivity of QPET to enable the support for more QoI formats in practical use cases; (2) Optimize the strategies for determining the point-wise and global error bounds in QoI-preserving error-bounded lossy compression; (3) Enhance the throughput of QPET to further reduce the computational cost for QoI preservation in the error-bounded lossy compression tasks;

### v4

@@ 6.2.1. Outperforming speeds
First, we measure and profile the execution speeds of baselines and QPET-integrated compressors to examine how QPET has accelerated the QoI-preserving error-bounded lossy compression. From the compression and decompression speeds on the Anvil cluster with the QoI functions $Q(x)=x^{2}$ and $Q(x)=x^{3}$ and the relative QoI error threshold $\tau=1\mathrm{e}{-3}$, reported in Figure 3 to 4 (no results from QoI-SZ3/HPEZ for $Q(x)=x^{3}$ because their current implementation don’t support this QoI), we can conclude as follows: compared to existing parameter-search-based baselines and QoI-preseving compressors, QPET-integrated compressors have significantly improved the compression throughputs (4x $\sim$ 6x of *-OptZ-R and 2x $\sim$ 2.5x of QoI-SZ3/HPEZ), with similar decompression throughput with the fast general-purpose compressors. Parameter search on general-purpose compressors (*-OptZ-R) needs to perform repeated executions for acquiring a certain QoI error threshold, and in our evaluations, this task may cost 4 to 10 or even more compression executions and QoI validations. Moreover, QPET brings quite limited time overheads in the decompression process (because neither error-bound computation nor QoI validation is needed), and sometimes it even improves the decompression speed over baselines due to the reduced accuracy required in the compression to ensure the same QoI error threshold. Last, compared to the existing QoI-SZ3/HPEZ, for the same use cases, QPET gains 100% $\sim$ 150% compression speed improvement and up to 400% decompression speed improvement, which is highly attributed to the efficient error-bound selection and tuning process proposed by QPET, and also the revised implementation in storing and recovering the point-wise error bounds.
@@ 6.2.2. Representative showcases
Table 5. Showcases of QoI-preserving error-bounded lossy compression. For blocked QoIs, $n=4^{3}$ (i.e., average on 4x4x4 blocks). $\epsilon$: data error bound. $\tau$: QoI error threshold. #It: Number of iterations. CR: Compression ratio. $T_{c}$: Compression throughput (in MB/s). $T_{d}$: Decompression throughput (in MB/s). All experiments are performed on Purdue Anvil, and all error bounds are relative (absolute value/value range). N/A indicates that the baseline is not available in the case.
@@
From those showcases, key findings are: (1) Matching the discussion in Section 6.2.1, QPET-integrated compressors deliver better compression throughputs than all baselines. The QPET-integrated compressor achieves 2x to 10x compression speedups over the parameter-search-based solutions. Moreover, SZ3/HPEZ-QPET has 2x to 3x compression speeds and 3x to 6x decompression speeds over QoI-SZ3/HPEZ. (2) With better throughputs, QPET-integrated compressors bring no worse (and mostly improved) compression ratios over the baselines. On $q(x)=x^{2}$, under error bounds of $10^{-4}$, QPET gained similar compression ratios with baselines in much shorter time costs. On preserving $q(x)=x^{3}$ of RTM-3500 and Hurricane-Cloud data (they both feature small absolute values less than 0.002), QPET-integrated compressors achieved 30% - 700% compression ratio improvements over the best parameter-search solution. On NYX-Velocity_x data ranging from zero to tens of millions, general-purpose compressors failed to preserve the QoI $\tanh{x}$ well due to the required point-wise data accuracy extremely diverging among small and large values. In contrast, QPET handled this case excellently. On multivariate QoIs, QPET significantly improved the compression ratio over baselines, as benefited from Theorem 5.4. Compared with existing QoI-preserving compressors on basic QoIs, QPET-integrated compressors still achieve noticeable compression ratio gains. On preserving $q(x)=\log_{2}x$ on Scale-LetKF-T data, SPERR-QPET achieved 100% CR improvement over QoI-HPEZ under $\epsilon=\tau=10^{-4}$. On preserving $q(X)=\frac{1}{n}\sum x_{i}^{2}$ on SegSalt-Pressure-3000 data, SPERR-QPET achieved 133% CR improvement over QoI-HPEZ under $\tau=10^{-3}$. General-purpose compressors usually over-preserve data quality to preserve the QoI error threshold because they can neither apply point-wise data accuracy nor auto-correct outliers. This explains why non-QoI-integrated compressors have limited compression ratios on QoI-preserving compression tasks, even after parameter searches on optimizing the compression ratio. Meanwhile, existing QoI-preserving compressors (like QoI-SZ3/HPEZ and MGARD-QoI) bring significant computational overheads for the QoI-preserving compression tasks. With lightweight error-bound tuning and outlier correction mechanisms, QPET successfully avoided the abovementioned limitations, proving its effectiveness and usability in preserving the QoI.
@@ 6.2.3. Improved Rate-distortions
To comprehensively understand the compression ratio improvement brought about by QPET in QoI-preserving error-bounded data compression, the compressors mentioned above are evaluated on a wide range of QoI functions and error thresholds. Then, we plot the bit rates and maximum QoI errors (relative, absolute value divided by value range), showing that QPET-integrated compressors have substantially improved the compression ratio when achieving the same maximum QoI errors as others. Due to the page limit, we will present the most representative results among the excellent outcomes we have achieved.
@@
Regarding different point-wise QoI functions, Figure 5 presents the bit rate-maximum QoI error plots obtained on 6 data domains. In this figure and all other similar ones in the paper, the results from the same base compressors are plotted as curves of the same color, with QPET-integrated compressors represented by solid lines and baselines by dashed lines. Moreover, the results from QoI-SZ3/HPEZ appear if and only if the corresponding) QoI is supported by them. In Figure 5, QPET-integrated compressors have shown the optimized compression ratio (bit rate) in every test case and achieve substantial improvements in most. For example, on the Scale-LetKF dataset (Figure 5 (b)), SPERR-QPET has around 100%/50% compression ratio improvements over the best baseline (QoI-SZ3/HPEZ) when the error threshold for the square of data ($Q(x)=x^{2}$) is set as $1\mathrm{e}{-2}$/$1\mathrm{e}{-3}$ (relative, and all the followings are same). On other QoIs for which existing QoI-preserving data compressors fail to support, such as $x^{3}$ and $\tanh{x}$ (Figure 5 (c, d, f)), QPET makes great use of the point-wise data accuracy to preserve QoIs, significantly outperforming parameter-search-based baselines (*-OptZ-R) by up to over 1000% compression ratio improvement (on NYX dataset with $Q(x)=x^{3}$) and higher throughputs. Regarding the preservation of $Q(x)=\sin{10x}$, which is close to a linear function in several local data regions, point-wise data accuracy contributes a limited amount to the compression. Therefore, the compression ratios of QPET-integrated compressors do not outperform parameter-search-based baselines. Nevertheless, QPET-integrated compressors can acquire the same compression ratios at much lower time costs, which is still a critical improvement for high-performance data processing tasks.
@@
Multivariate QoI functions contain two sub-categories: QoI on data blocks within one data field and QoI vector data among multiple fields. In Figure 6, for better comparison with baselines (particularly QoI-SZ3/HPEZ and MGARD-QoI), we perform evaluations of QPET and the baselines on the average of $x$ (data value itself), $x^{2}$ (square of the data value), and $x^{3}$ (cubic of the data value) on partitioned fixed-size data blocks. In Figure 6 (a) and (b), we can find that MGARD-QoI and QoI-SZ3 present unstable and suboptimal compression ratios, and QoI-HPEZ cannot outperform HPEZ-QPET and SPERR-QPET with an obvious gap between them. For $Q(X)=\frac{1}{n_{b}}\sum x^{2}$ and $Q(X)=\frac{1}{n_{b}}\sum x^{3}$, the 3 QPET-integrated compressors are the top-3 of all compressors, outperforming both parameter-search-based baselines and QoI-SZ3/HPEZ. On preserving the $Q(X)=\frac{1}{n_{b}}\sum x^{2}$ when compressing the RTM dataset (Figure 6 (c)), SPERR-QPET gains an aggregated compression ratio of $\approx$ 215 (bit rate of $\approx$0.15) under the QoI error tolerance of $1\mathrm{e}{-4}$, which is 108% better than the best baseline SPERR-OptZ-R (CR $\approx$ 103). On preserving the $Q(X)=\frac{1}{n_{b}}\sum x^{3}$ when compressing the Hurricane dataset (Figure 6 (e)), HPEZ-QPET gains aggregated compression ratios of $\approx$ 121/45.4 under the QoI error tolerance of $1\mathrm{e}{-3}$/$1\mathrm{e}{-4}$, which is 200%/220% of HPEZ-OptZ-R compression ratios ( $\approx$ 60.9/20.6). The improvement of compression ratios in those test cases is highly attributed to Theorem 5.4 because of the large number of variables (same as the block size 64) in the QoIs so that QPET can apply much higher point-wise QoI ($x$/$x^{2}$/$x^{3}$) error thresholds on each data point compared to the overall error threshold.
@@ 6.2.4. Ablation study
At the end of our evaluation, we conduct an ablation study of QPET design components, separately validating how different strategies described in Algorithm 2,  3, and  4 contribute to the performance of QPET.
On bounding the error of $Q(x)=x^{3}$, Figure 8 (a) and (b) compare QPET to its derivation QPET-A, which uses the analytical solution of the inequality $\left|(\epsilon_{i}+x_{i})^{3}-x_{i}^{3}\right|\leq\tau$ instead of Algorithm 2 to determine the point-wise error bound $\epsilon_{0}$ for $x_{0}$. On both datasets, QPET achieves the same compression ratio as QPET-A, indicating that Algorithm 2 performs well in replacing analytic solutions. This fact indicates that the mechanism of QPET has worked well in determining the compression error bounds.
Next, we investigate the impact of Throrem 5.4 and the parameter $c$ by evaluating SPERR-QPET on the SCALE-LetKF with the QoI of the average of $x^{2}$ on 4$\times$4$\times$4 data blocks with different configurations. Under $c=0$, SPERR-QPET deactivates Theorem 5.4, and larger $c$ brings large error bound estimations from Throrem 5.4. Figure 8 (c) illustrates the corresponding results. Higher $c$ does improve compression by proposing higher and still trustworthy estimations of point-wise error bounds. As the improvement in compression ratios becomes minor when $c$ reaches 3.0, we selected this value as the default in SPERR-QPET.
@@ 7. Conclusion
To meet the significant requirements of Quantity-of-Interest (QoI) preserving in error-bounded lossy compression and address the critical limitations of existing solutions, we propose QPET, a versatile and portable framework for error-bounded lossy compression that enables the preservation of a broad range of QoI while significantly improving the performance and quality of existing state-of-the-arts. We integrate QPET into 3 representative error-bounded lossy compressors of different archetypes. Experimental results demonstrate that QPET is faster than all existing QoI-preserving error-bounded lossy compression solutions, leading to 2x to 10x compression speedups over the baselines of both parameter-search methods and direct QoI-preserving methods. QPET also delivers up to 1000% compression ratio improvements over existing QoI-preserving error-bounded lossy compression solutions. In future research, we will optimize QPET in the following three aspects: (1) Enable the support for more QoI formats in practical use cases; (2) Optimize the strategies for determining the point-wise and global error bounds in QoI-preserving; (3) Enhance the throughput to further reduce the computational cost for QoI preservation;
8. Acknowledgments
This research was supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research (ASCR), under contracts DE-AC02-06CH11357. This work was also supported by the National Science Foundation (Grant Nos. 2104023, 2311875, 2344717, OAC-2330367, OAC-2313122, OAC-2311756). We also acknowledge the computing resources provided by Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) – Purdue Anvil.